Find the areas of the regions enclosed by the lines and curves y = 2vx and y = 2x - 4 20 7 19 B) 6 D) 19 26 1 F) H) 3 3 2 NIMchoose correct answer
Answer #1

To find the area enclosed by: =255 and y = 2800 y= 22-4 y² = 40 4 -2012 y = 2 y² (4,4) Required Area = 0 dy 4 (9,0) y+4 En 2